# A Topological Gaussian Mixture Model for Bone Marrow Morphology in Leukaemia Qiquan Wang1,†, Anna Song1,2, Antoniana Batsivari2, Dominique Bonnet2, and Anthea Monod1

### v1

@@ Abstract
Acute myeloid leukaemia (AML) is a type of blood and bone marrow cancer characterized by the proliferation of abnormal clonal haematopoietic cells in the bone marrow leading to bone marrow failure. Over the course of the disease, angiogenic factors released by leukaemic cells drastically alter the bone marrow vascular niches resulting in observable structural abnormalities. We use a technique from topological data analysis—persistent homology—to quantify the images and infer on the disease through the imaged morphological features. We find that persistent homology uncovers succinct dissimilarities between the control, early, and late stages of AML development. We then integrate persistent homology into stage-dependent Gaussian mixture models for the first time, proposing a new class of models which are applicable to persistent homology summaries and able to both infer patterns in morphological changes between different stages of progression as well as provide a basis for prediction.
@@ 1 Introduction
Acute myeloid leukaemia (AML) is a cancer of the bone marrow and blood which can also infiltrate other haematopoietic organs such as the spleen and other tissues (Behrmann et al.,, 2018). It is characterized by the proliferation of abnormally differentiated and nonfunctional haematopoietic blast cells, which hijack and alter the bone marrow niches; they suppress normal haematopoiesis in favor of their own growth (Kumar et al.,, 2017). AML cells are known to intervene with the normal microenvironment equilibrium and remodel the bone marrow (Passaro et al.,, 2017; Duarte et al.,, 2018), giving rise to prominent morphological changes such as increases in the vessel density and vessel leakiness. Overall, the regularity in the bone marrow vasculature degrades, leading to poorer fluid circulation which in turn further exacerbates the disease. Data on AML are difficult and very costly to obtain and analyze, which contribute to the challenging nature of the disease study.
In this paper, we both quantify and model the progressive alteration of the vessels in the bone marrow microenvironment, captured using 3D confocal microscopy, as AML develops. We utilize a central methodology in topological data analysis—persistent homology—to quantitatively summarize both the local and global the geometric information within the imaging data in an intepretable manner, as scatter plots in the plane known as persistence diagrams. This then allows the geometric information to be modeled from the observed patterns and clusters in the persistence diagrams. We then propose and implement a new class of stage-dependent Gaussian mixture models to model the deformation of the bone marrow architecture within the progression of AML across its stages.
@@
Related Work.
Our understanding of the morphological progression is far from complete, despite numerous prior studies. Among those using images to understand the disease, as is our aim in this work, Gomariz et al., (2018); Coutu et al., (2017, 2018) quantifiably analyzed the spatial distribution and interactions of different cells and features within the bone marrow niches in high-resolution microscopy images. Shih et al., (2009); Gomes et al., (2021) conducted similar analyses on the leukaemic-induced changes using magnetic resonance images.
Persistent homology has been successfully implemented to study the shapes within data in medical imaging in oncology. Notable examples include classifying hepatic tumors from magnetic resonance images (Oyama et al.,, 2019); tumor segmentation from whole slide images (Qaiser et al.,, 2016); computer-aided diagnosis (Yu et al.,, 2016); and prognosis (Crawford et al.,, 2020). On a finer level, persistent homology has also been used to summarize cellular architecture in applications to breast cancer (Singh et al.,, 2014; Aukerman et al.,, 2022) and prostate cancer (Lawson et al.,, 2019).
@@ 2 Data Description: Confocal Imaging
We study the morphological changes to bone marrow vascular niches caused by AML under control environments by engrafting AML and non-AML cells into immunodeficient mice. AML, human primary samples, and leukaemic cell lines of different types and clonalities were injected intravenously, and the mice were sacrificed at different stages as the disease developed. From each mouse, we retained one of the femur bones for immunostaining and 3D confocal imaging, while other bones were crushed and used to approximate the percentage of engraftment. These data are very costly to obtain in terms of time consumption as well as equipment and resources required.
@@
When AML cells are engrafted into the bone marrow, they release cytokines—growth factors and other molecules which alter the bone marrow niche into one which favors the proliferation of AML cells. Observable alterations include increases in the vessel density, proportion of endothelial cells, and vascular leakiness, as well as vessel compression, hypoxia, and altered perfusion (Passaro et al.,, 2017). Figure 2 shows a comparison between the vessel structure in corresponding regions in a healthy bone marrow and a bone marrow engrafted with AML. We observe that within the healthy bone marrow, the vessels form a smooth, homogeneous, and regular network, whereas in the bone marrow after engraftment, the vessels become thinner, inhomogeneous, more permeable, and irregular.
@@
6 P1 samples: Samples injected with “patient 1’s” AML cells at 10%, 40%, 44%, 51%, 60%, 76% engraftment (“patient 1” is a pseudo ID).
3 P2 samples: Samples injected with “patient 2’s” AML cells at 59%, 88%, 90% engraftment (“patient 2” is a pseudo ID).
Biologically, we expect the CTRL samples to display normal murine vessel structure within the bone marrow since these samples were not subjected to any alterations. The MNC cells were harvested from blood sampled from the umbilical cord of human newborns, consisting of a variety of blood cells with a single nucleus, such as lymphocytes, monocytes, eosinophils, basophils, and various stem/progenitor cells. These samples are designed to play the role of “fake” control samples for when healthy human cells are engrafted into the murine bone marrow. U937 and HL60 bone samples were injected with different types of cultured immortalized cell lines from human extracts, meaning that the injected population of cells had particular mutations which allowed them to evade normal cellular senescence and proliferate indefinitely. These cells were cultured to retain certain interesting phenotypes and functions before being injected into mice. Notably, these cell lines are monoclonal, meaning that the whole population of cells descends from the same ancestor cell. In contrast, P1 and P2 bone samples were injected with polyclonal cells extracted from human patients, which means that the collections of cells have much more complex mutation landscapes and there is no control over the types of clonal cells that were injected. There are several possible scenarios for the engraftment of these cells in the bone marrow. For example, some clonal subpopulations may be more aggressive and grow more rapidly than others. Alternatively, different clonal subpopulations may cohabitate without singular or group dominance in the whole population. The spatial distribution of these subpopulations is unknown.
Preprocessing and Image Segmentation.
Confocal microscopy of the bone samples generates multiple 2D slices along the depth of the samples which can be stacked together to recreate the 3D images. To preprocess the stack of confocal microscopy images embedded with noise and distortions, we pass a median filter and apply Niblack local thresholding (Niblack,, 1985) to the images to approximate the positions of the vessels and cavities in the image slices. We reconstruct the 3D structures, ensuring linkage between the bone marrow vasculature within each slice using a Willmore regularization method (Bretin et al.,, 2017; Song,, 2022). Details of the image segmentation and reconstruction process can be found in Song, (2023).
@@ 3.1 Persistent Homology
Persistent homology is a methodology that adapts the concept of homology from algebraic topology to the computational setting, notably for real data analysis. Classical homology algebraically counts the number and types of holes of sets and spaces, as a means to characterize them and distinguish between sets and spaces. Here, “holes” correspond to connected components, loops, voids, and higher dimensional analogues; specifically, we are interested in those holes which are invariant to continuous deformations, such as stretching and compressing. The degree of homology corresponds to different types of holes: degree 0 homology captures the connected components; degree 1 homology captures the loops; degree 2 homology captures the voids or “bubbles,” and so on. In the context of our application, we are interested in the first 3 degrees of homology, where intuitively, degree 0 homology should provide information corresponding to the network of interconnected vessels running throughout the bone, while degrees 1 and 2 homology capture loops and empty nodules created within such a network.
@@
Computing homology is a difficult task, so computational topology techniques largely rely on simplicial homology, where the set or space of interest is discretized into a skeletal structure, where there exist tractable and efficient algorithms to compute homology. The discretization of the set or space of interest represents it as a simplicial complex, which is composed of simplices—vertices, edges, triangles, and higher dimensional structures—assembled in a combinatorial manner, where every face of a simplex in the complex must also belong to that complex. We can then compute the homology of the simplicial complex combinatorially; we use to denote the th homology group consisting of -simplices.
@@
In topological data analysis, the approach to discretization of our set or space is important: typically, this is done by sampling a number of points from the set or space, then connecting them to create the simplicial complex by a predetermined connection rule. A popular and intuitive choice for computation is the Vietoris–Rips complex, where the connection rule is that two points are joined by an edge if two balls with the points as centers overlap; higher dimensional structures are created in a similar manner by considering the overlapping of the balls for more points. A crucial question in topological data analysis, then, is to determine the “correct” radius of the balls in order to construct the simplicial complex to then compute homology. The answer that persistent homology gives is to consider a spectrum of radii, rather than one fixed radius, in order to produce a multiresolutional summary of the homology. The growth of the radii are dictated by a continuous, real-valued filtration function, which produces a filtration—an increasing collection of nested subcomplexes to the simplicial complex representing the discretized set or space of interest—according to the predetermined connection rule.
@@
for . This is referred to as the sublevel set filtration, with the property that for all . An illustration of a sublevel set filtration dictated by the height function is given in Figure 4, originally appearing in Crawford et al., (2020).
Persistent homology computes the homology over this sequence of subcomplexes and allows us to track how the topological features of the data, summarized by the number and types of simplices, evolve over the different resolutions. Each feature is tracked by an interval, where the left endpoint (usually referred to as the birth time) indicates when the feature appeared, and the right endpoint (usually referred to as the death time) indicates when that feature disappears (by absorption into a larger structure), lending interpretability to the procedure: longer intervals usually refer to “more important” features while shorter intervals are usually considered to be topological “noise.” Formally, this interpretability is given by the structure theorem for persistent homology (Crawley-Boevey,, 2014), where the geometric and topological information encoded by the homology groups together with the mapping induced by the inclusions between the sublevel sets generate a persistence module for a fixed degree , which decomposes uniquely (up to isomoprhism) into a direct sum of irreducible interval modules. The collection of these intervals is the output of persistent homology; it is typically plotted as a scatterplot in the upper half plane, together with the diagonal (i.e., the set of zero-length intervals where birth is equal to death), in a persistence diagram.
@@ 3.2 Signed Distance Persistent Homology
Song et al., (2023) develop a generalized Morse theory of the signed distance function in the context of persistent homology; Morse theory is a branch of differential topology that studies topological features using differentiable functions, specifically, in terms of critical points. By Corollary 5 of Song et al., (2023), for a generically embedded surface in , the signed distance function is a topological Morse function with finitely many critical points. In this setting, there exists a direct correspondence between the non-degenerate critical values of the signed distance function and the interval decomposition of the corresponding persistence module for fixed degree , in particular, a correspondence to either the birth or death times of topological features captured by homology.
@@ 4 Quantifying AML with Signed Distance Persistent Homology
In this section, we present our first contributions, which is a quantification and analysis of AML bone marrow vasculature imaging using SDPH. In particular, we show that SDPH is able to rigorously capture the quantitative information within confocal bone marrow vasculature imaging in a way that is interpretable and indicative of different disease stages, given the data (see Section 2).
@@ 4.1 Applying SDPH and Data Interpretation
Song et al., (2023) show that there are six types of non-degenerate critical points of the signed distance function, each with its individual geometric interpretation and corresponding to either a birth or death time of a homological component. In their original notation (see Song et al., (2023)), the six types of critical points were referred to as the , , , , , and critical points, corresponding to their Morse index and whether the points occur within or outside the surface. In this paper for convenience and ease of interpretation, we will alter the notation and rename the critical points as the , , , , , and sizes, respectively, so that they reflect the absolute values in the original notation. This relabeling is done with the intuition that the , , and sizes measure different behaviors in radius or thickness of vessels, and that the , , and measure behaviors in the in interstitial gaps between vessels in our biomedical imaging application. With this relabeled notation, the morphological interpretation of the types of critical points as measured inside the bone marrow architecture are given as follows.
@@
Figures from Song et al., (2023) illustrate the local behavior found near different types of critical points; a selection is presented in Figure 5. In these illustrations, the red dots represent the non-degenerate critical points and the black dots their closest point(s) on the surface, respectively. In Figure 5(a), we observe the behavior around either a or critical point. The difference arises from whether we consider the critical point to be inside or outside the lumen. If the point is inside the lumen, then the construction mimics a critical point of type found within a biconcavity of a dimple on a blood vessel. Alternatively, if the point is outside the lumen, then the two pieces of surface in the construction represent the outside linings of two vessels and a critical point of type found between them in the narrowing formed. Both Figures 5(b) and 5(c) illustrate the possible local morphologies around a type or critical point. In Figure 5(b), the critical point is found within two concave pieces of surface, which may either be the vessel walls around a critical point or the inner lumen of a concavity surrounded by vessels. Figure 5(c) more clearly shows the construction of a critical point found within an open cavity.
@@
The pairings of critical points within different quadrants present further geometric interpretations. For example, within the PH1 north-west (NW) quadrant, the number of birth–death points corresponds to the number of loops in the sample, while the individual birth and death times measure the thicknesses and sizes of the loops up to a difference in sign. We also note that the pairing of and critical sizes should rarely occur for vessels as this alludes to a bubble of extravascular space fully trapped inside the lumen, which intuitively should not normally appear.
@@
We utilize the critical sizes in analysis of the bone marrow imaging data in two different ways: one is a “local” texture analysis and the other is a “global” texture analysis. In the local analysis, we compute the morphology (i.e., the SDPH diagrams and its summary of critical points) over each image segmented into numerous small regions with the goal of distilling the heterogeneous textures into mixtures of similar and more homogeneous textures through clustering. Then, based on this decomposition into homogeneous textures for each image, we further group the images into what we refer to as phases, which correlate with the level of engraftment. For the global analysis, we instead compute the morphology over the whole images and using the density of points on the SDPH diagrams, we further categorize the image samples into groups which are in concordance with the results of the local analysis.
It is important to note that there are distinctions between the following terms used in this paper: The actual stage of engraftment of the AML or injected cells within the mice, which varies from region to region and cannot be directly measured, which differs from the estimated stage of engraftment, which is a global estimate of the amount of engraftment through the process involving crushing all the remaining bones of the mice other than imaged femur bone imaged. We refer to the estimated stage of engraftment in this paper as simply the level of engraftment. We refer to clusters of similar morphology evaluated with local analysis as textures and the overall spread of different morphology over the bone as mixture of textures. We assign an arbitrary labeling to these textures: A, B, and C.

### v2

# A Topological Gaussian Mixture Model for Bone Marrow Morphology in Leukaemia Qiquan Wang1,ā , Anna Song1,2, Antoniana Batsivari2, Dominique Bonnet2, and Anthea Monod1

@@ Abstract
Acute myeloid leukaemia (AML) is a type of blood and bone marrow cancer characterized by the proliferation of abnormal clonal haematopoietic cells in the bone marrow leading to bone marrow failure. Over the course of the disease, angiogenic factors released by leukaemic cells drastically alter the bone marrow vascular niches resulting in observable structural abnormalities. We use a technique from topological data analysisāpersistent homologyāto quantify the images and infer on the disease through the imaged morphological features on a proprietary high-resolution image dataset. We find that persistent homology uncovers succinct dissimilarities between the control, early, and late stages of AML development. We then integrate persistent homology into stage-dependent Gaussian mixture models, proposing a new class of models which are applicable to persistent homology summaries and able to both infer patterns in morphological changes between different stages of progression as well as provide a basis for prediction.
@@ 1 Introduction
Acute myeloid leukaemia (AML) is a cancer of the bone marrow and blood which can also infiltrate other haematopoietic organs such as the spleen and other tissues (Behrmann etĀ al.,, 2018). It is characterized by the proliferation of abnormally differentiated and nonfunctional haematopoietic blast cells, which hijack and alter the bone marrow niches; they suppress normal haematopoiesis in favor of their own growth (Kumar etĀ al.,, 2017). AML cells are known to intervene with the normal microenvironment equilibrium and remodel the bone marrow (Passaro etĀ al.,, 2017; Duarte etĀ al.,, 2018), giving rise to prominent morphological changes such as increases in the vessel density and vessel leakiness. Overall, the regularity in the bone marrow vasculature degrades, leading to poorer fluid circulation which in turn further exacerbates the disease. Data on AML are difficult and very costly to obtain and analyze, which contributes to the challenging nature of the disease study.
In this paper, we both quantify and model the progressive alteration of the vessels in the bone marrow microenvironment, captured using 3D confocal microscopy at high resolutions in a proprietary dataset, as AML develops. We utilize a central methodology in topological data analysisāpersistent homologyāto quantitatively summarize both the local and global geometric information within the imaging data in an interpretable manner, as scatter plots in the plane known as persistence diagrams. This allows the geometric information to be modeled from the observed patterns and clusters in the persistence diagrams. We then propose and implement a new class of stage-dependent Gaussian mixture models to model the deformation of the bone marrow architecture with the progression of AML across its stages.
@@
Related Work. Current understanding of the morphological progression is far from complete, despite numerous prior studies. Among those using images to understand the disease, as is our aim in this work, Gomariz etĀ al., (2018); Coutu etĀ al., (2017, 2018) quantifiably analyzed the spatial distribution and interactions of different cells and features within the bone marrow niches in high-resolution microscopy images. Shih etĀ al., (2009); Gomes etĀ al., (2021) conducted similar analyses on the leukaemic-induced changes using magnetic resonance images. More broadly, a recent survey by Mohammed etĀ al., (2024) summarized some of the modern techniques for quantitative cancer imaging analysis.
Persistent homology has been successfully implemented to study the shapes within medical imaging data in oncology. Notable examples include the classification of hepatic tumors from magnetic resonance images (Oyama etĀ al.,, 2019); classification of acute lymphoblastic leukaemia from microscopy images Shah etĀ al., (2024); tumor segmentation from whole slide images (Qaiser etĀ al.,, 2016); computer-aided diagnosis (Yu etĀ al.,, 2016); and prognosis (Crawford etĀ al.,, 2020). More recent work by Pritchard etĀ al., (2023) and Moon etĀ al., (2020) demonstrated the use of signed Euclidean distance transform on images to quantify bone microstructure and tumor morphology for classification and prognostics. On a finer level, persistent homology has also been used to summarize cellular architecture in applications to breast cancer (Singh etĀ al.,, 2014; Aukerman etĀ al.,, 2022) and prostate cancer (Lawson etĀ al.,, 2019).
@@ 2 Data Description: Confocal Imaging
We study the morphological changes to bone marrow vascular niches caused by AML under control environments by engrafting AML and non-AML cells into immunodeficient mice. AML, human primary samples, and leukaemic cell lines of different types and clonalities were injected intravenously, and the mice were sacrificed at different stages as the disease developed. From each mouse, we retained one of the femur bones for immunostaining and 3D confocal imaging, while other bones were crushed and used to approximate the percentage of engraftment. These data are very costly to obtain in terms of time consumption as well as equipment and resources required: they are among the first to study the structure of the bone marrow vascular niches at such a fine resolution, crucial for the novel investigation into differences between early and late stages of AML as most research has focused on identifying changes and causes in late stages.
@@
When AML cells are engrafted into the bone marrow, they release cytokinesāgrowth factors and other molecules which alter the bone marrow niche into one which favors the proliferation of AML cells. Observable alterations include increases in vessel density, proportion of endothelial cells, vascular leakiness, as well as vessel compression, hypoxia, and altered perfusion (Passaro etĀ al.,, 2017). Figure 2 shows a comparison between the vessel structure over corresponding regions in a healthy bone marrow and a bone marrow engrafted with AML. We observe that within the healthy bone marrow, the vessels form a smooth, homogeneous, and regular network, whereas in the bone marrow after engraftment, the vessels become thinner, inhomogeneous, more permeable, and irregular.
@@
ā¢
6 P1 samples: Samples injected with āpatient 1āsā AML cells at 10%, 40%, 44%, 51%, 60%, 76% engraftment (āpatient 1ā is a pseudo ID).
ā¢
3 P2 samples: Samples injected with āpatient 2āsā AML cells at 59%, 88%, 90% engraftment (āpatient 2ā is a pseudo ID).
Biologically, we expect the CTRL samples to display normal murine vessel structure within the bone marrow since these samples were not subjected to any alterations. The MNC cells were harvested from blood sampled from the umbilical cord of human newborns, consisting of a variety of blood cells with a single nucleus, such as lymphocytes, monocytes, eosinophils, basophils, and various stem/progenitor cells. These samples are designed to play the role of āfakeā control samples for when healthy human cells are engrafted into the murine bone marrow. U937 and HL60 bone samples were injected with different types of cultured immortalized cell lines from human extracts, meaning that the injected population of cells had particular mutations which allowed them to evade normal cellular senescence and proliferate indefinitely. These cells were cultured to retain certain interesting phenotypes and functions before being injected into mice. Notably, these cell lines are monoclonal, meaning that the whole population of cells descends from the same ancestor cell. In contrast, P1 and P2 bone samples were injected with polyclonal cells extracted from human patients, which means that the collections of cells have much more complex mutation landscapes and there is no control over the types of clonal cells that were injected. There are several possible scenarios for the engraftment of these cells in the bone marrow. For example, some clonal subpopulations may be more aggressive and grow more rapidly than others. Alternatively, different clonal subpopulations may cohabitate without singular or group dominance in the whole population. The spatial distribution of these subpopulations is unknown.
Preprocessing and Image Segmentation. Confocal microscopy of the bone samples generates multiple 2D slices along the depth of the samples which can be stacked together to recreate the 3D images. To preprocess the stack of confocal microscopy images embedded with noise and distortions, we pass a median filter and apply Niblack local thresholding (Niblack,, 1985) to the images to approximate the positions of the vessels and cavities in the image slices. We reconstruct the 3D structures, ensuring linkage between the bone marrow vasculature within each slice using a Willmore regularization method (Bretin etĀ al.,, 2017; Song,, 2022). Details of the image segmentation and reconstruction process can be found in Song, (2023).
@@ 3.1 Persistent Homology
Persistent homology is a methodology that adapts the concept of homology from algebraic topology to the computational setting, notably for real data analysis. Classical homology algebraically counts the number and types of holes of sets and spaces, as a means to characterize them and distinguish between sets and spaces. Here, āholesā correspond to connected components, loops, voids, and higher dimensional analogues; specifically, we are interested in those holes which are invariant to continuous deformations, such as stretching and compressing. The degree of homology corresponds to different types of holes: degree 0 homology captures the connected components; degree 1 homology captures the loops; degree 2 homology captures the voids or ābubblesā, and so on. In the context of our application, we are interested in the first 3 degrees of homology, where intuitively, degree 0 homology should provide information corresponding to the network of interconnected vessels running throughout the bone, while degrees 1 and 2 homology capture loops and empty nodules created within such a network.
@@
Computing homology is a difficult task, so computational topology techniques largely rely on simplicial homology, where the set or space of interest is discretized into a skeletal structure, where there exist tractable and efficient algorithms to compute homology. The discretization of the set or space of interest represents it as a simplicial complex, which is composed of simplicesāvertices, edges, triangles, and higher dimensional structuresāassembled in a combinatorial manner, where every face of a simplex in the complex must also belong to that complex. We can then compute the homology of the simplicial complex combinatorially; we use to denote the th homology group consisting of -simplices.
@@
In topological data analysis, the approach to discretization of our set or space is important: typically, this is done by sampling a number of points from the set or space, then connecting them to create the simplicial complex by a predetermined connection rule. A popular and intuitive choice for computation is the VietorisāRips complex, where the connection rule is that two points are joined by an edge if two balls with the points as centers overlap; higher dimensional structures are created in a similar manner by considering the overlapping of the balls for more points. A crucial question in topological data analysis, then, is to determine the ācorrectā radius of the balls in order to construct the simplicial complex to then compute homology. The answer that persistent homology gives is to consider a spectrum of radii, rather than one fixed radius, in order to produce a multiresolutional summary of the homology. The growth of the radii are dictated by a continuous, real-valued filtration function, which produces a filtrationāan increasing collection of nested subcomplexes to the simplicial complex representing the discretized set or space of interestāaccording to the predetermined connection rule.
@@
for . This is referred to as the sublevel set filtration, with the property that for all . An illustration of a sublevel set filtration dictated by the height function is given in Figure 4, originally appearing in Crawford etĀ al., (2020).
Persistent homology computes the homology over this sequence of subcomplexes and allows us to track how the topological features of the data, summarized by the number and types of simplices, evolve over the different resolutions. Each feature is tracked by an interval, where the left endpoint (usually referred to as the birth time) indicates when the feature appeared, and the right endpoint (usually referred to as the death time) indicates when that feature disappears (by absorption into a larger structure), lending interpretability to the procedure: longer intervals usually refer to āmore importantā features while shorter intervals are usually considered to be topological ānoise.ā Formally, this interpretability is given by the structure theorem for persistent homology (Crawley-Boevey,, 2014), where the geometric and topological information encoded by the homology groups together with the mapping induced by the inclusions between the sublevel sets generate a persistence module for a fixed degree , which decomposes uniquely (up to isomoprhism) into a direct sum of irreducible interval modules. The collection of these intervals is the output of persistent homology; it is typically plotted as a scatterplot in the upper half plane, together with the diagonal (i.e., the set of zero-length intervals where birth is equal to death), in a persistence diagram.
@@ 3.2 Signed Distance Persistent Homology
Song etĀ al., (2025) developed a generalized Morse theory of the signed distance function in the context of persistent homology; Morse theory is a branch of differential topology that studies topological features using differentiable functions, specifically, in terms of critical points. By Corollary 5 of Song etĀ al., (2025), for a generically embedded surface in , the signed distance function is a topological Morse function with finitely many critical points. In this setting, there exists a direct correspondence between the non-degenerate critical values of the signed distance function and the interval decomposition of the corresponding persistence module for fixed degree , in particular, a correspondence to either the birth or death times of topological features captured by homology.
@@ 4 Quantifying AML with Signed Distance Persistent Homology
In this section, we present our first contributions, which is a quantification and analysis of AML bone marrow vasculature imaging using SDPH. In particular, we show that SDPH is able to rigorously capture the quantitative information within confocal bone marrow vasculature imaging in a way that is interpretable and indicative of different disease stages, given the data.
@@ 4.1 Applying SDPH and Data Interpretation
Song etĀ al., (2025) show that there are six types of non-degenerate critical points of the signed distance function, each with its individual geometric interpretation and corresponding to either a birth or death time of a homological component. In their original notation (see Song etĀ al., (2025)), the six types of critical points were referred to as the , , , , , and critical points, corresponding to their Morse index and whether the points occur within or outside the surface. In this paper for convenience and ease of interpretation, we will alter the notation and rename the critical points as the , , , , , and sizes, respectively, denoting the absolute values in the original notation. This relabeling is done with the intuition that the , , and sizes measure different behaviors in radius or thickness of vessels, and that the , , and measure behaviors in the in interstitial gaps between vessels in our imaging application. With this relabeled notation, the morphological interpretation of the types of critical points as measured inside the bone marrow architecture are given as follows.
ā¢
@@
Figures from Song etĀ al., (2025) illustrate the local behavior found near different types of critical points; a selection is presented in Figure 5. In these illustrations, the red dots represent the non-degenerate critical points and the black dots their closest points on the surrounding surfaces respectively. In Figure 5(a), we observe the behavior around either a or critical point. The difference arises from whether we consider the critical point to be inside or outside the lumen. If the point is inside the lumen, then the construction mimics a critical point of type found within a biconcavity of a dimple on a blood vessel. Alternatively, if the point is outside the lumen, then the two pieces of surface in the construction represent the outside linings of two vessels and a critical point of type found between them in the narrowing formed. Both Figures 5(b) and 5(c) illustrate the possible local morphologies around a type or critical point. In Figure 5(b), the critical point is found within two concave pieces of surface, which may either be the vessel walls around a critical point or the inner lumen of a concavity surrounded by vessels. Figure 5(c) more clearly shows the construction of a critical point found within an open cavity.
@@
The pairings of critical points within different quadrants present further geometric interpretations. For example, within the PH1 north-west (NW) quadrant, the number of birthādeath points corresponds to the number of loops in the sample, while the individual birth and death times measure the thicknesses and sizes of the loops up to a difference in sign. We also note that the pairing of and critical sizes should rarely occur for vessels as this alludes to a bubble of extravascular space fully trapped inside the lumen, which intuitively should not normally appear.
@@
We utilize the critical sizes in analysis of the bone marrow imaging data in two different ways: one is a ālocalā texture analysis and the other is a āglobalā texture analysis. In the local analysis, we compute the morphology (i.e., the SDPH diagrams and its summary of critical points) over each image segmented into numerous small regions with the goal of distilling the heterogeneous textures into mixtures of similar and more homogeneous textures through clustering. Then, based on this decomposition into homogeneous textures for each image, we further group the images into what we refer to as phases, which correlate with the level of engraftment. For the global analysis, we instead compute the morphology over the whole images and using the density of points on the SDPH diagrams, we find that the clustering behaviors of the approximate densities are in concordance with the results of the local analysis.
It is important to note that there are distinctions between the following terms used in this paper: The actual stage of engraftment of the AML or injected cells within the mice, which varies from region to region and cannot be directly measured, which differs from the estimated stage of engraftment, which is a global estimate of the amount of engraftment through the process involving crushing all the remaining bones of the mice other than imaged femur bone imaged. We refer to the estimated stage of engraftment in this paper as simply the level of engraftment. We refer to clusters of similar morphology evaluated with local analysis as textures and we observe the imaged samples as being composed of different textures.